if we consider that 1/6 in place of 1/12 mass of carbon atom is taken to be the relative atomic mass unit the mass of one mole of substance WILL.

Grade:12th pass

2 Answers

Neeti
8 years ago
remain unchanged.
 
This is because when you consider 1/6 in place of 1/12, since mass of carbon atom is still the same, new amu1 = 2amu ( where amu1 is new amu unit and amu is old amu unit)
 
So now Na1 (avogadro’s number) becomes half ( Na/2) since the weight of carbon corresponding to atomic weight becomes 6gm in place of 12.( where Na1 is new avogardo’s number and Na is old avogardro’s number)
 
And we know that mass of one mole of substance is = Na x amu
 
So new atomic mass= Na1 x amu1 = Na/2 x 2amu = Na x amu = old atomic weight.
 
 
Please feel free to ask any follow up doubt and please approve the answer if you think it’s useful :)
Yash Chourasiya
3 years ago
Hello Student

Mass of 1 mole substance =Na (Avogadro's constant) × mass of 1 atom (amu).
1 amu=(1/12)× mass of 1 atom of C-12.
After we assume it to be (1/6),1amu gets doubled.
Now,Na (Avogadro's constant) becomes half (Na/2) because the weight of carbon corresponding to atomic weight is now 6gm in place of 12gm.
As mass of 1 mole substance = Na×1amu.
For our new system,we get mass of 1 mole substance = (Na/2) × (2amu) = Na×1amu, giving us the initial result.
Hence, we can conclude that even by assuming 1/6 in place of 1/12 as the relative atomic mass unit, we get the same mass in 1 mole substance.

So it will remain unchanged.

I hope this answer will help you.